OpenStudy (anonymous):

At a certain time of the day, a 4-meter-tall vertical pole casts a shadow of 3 meters. What is the angle of elevation, to the nearest degree, of the Sun?

OpenStudy (anonymous):

tanθ = 4/3 θ = arctan(4/3) = 53.13 θ = 53°